OpenStudy (anonymous):

A direct variation equation is graphed. The point (6, 4) is on the graph. Which of the following points are also on the graph of the equation? (3, 2) (-9, -6) (-12, 8) (-6, -3) ((select all that apply)

OpenStudy (anonymous):

@Cookie_2046 do you know how to do this

OpenStudy (cookie_2046):

Not really

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

@kelliegirl33 do you know how to do this

OpenStudy (anonymous):

I tend to do direct variation a little bit differently....I use proportions...and proportions are nothing but equivalent fractions. (6,4)...x = 6 and y = 4 lets compare it to (3,2) 6/4 = 3/2 (cross multiply) (6)(2) = (3)(4) 12 = 12 (correct)....this one is an answer choice (6,4) and (-9,-6) 6/4 = -9/-6 (cross multiply) (6)(-6) = (4)(-9) -36 = -36 (correct).....this one is also an answer choice do you see how I am doing this ?? if the proportions do not come out equal, then it is not an answer choice. This way can only be done with direct variations...not indirect.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

so would (12,8) also be an answer because when you cross multiply, it comes out as 48 ?

OpenStudy (anonymous):

but your answer choice is (-12,8)....and that is not on the line

OpenStudy (anonymous):

oh oops haha i didnt see the "-"